# Parameterized sequential functions for temporal properties

### v2

@@ 5 Discussion
When Professor Krithivasan Ramamritham gave me temporal logic papers to read[12] at the start of my graduate studies in the 1980s, I had just finished working on a commerical distributed operating system[2] and was frustrated by how often the developers had found design errors only during test (after writing enormous quantities of code). Temporal logic seemed like a solution, but I was not unique in finding that it was not a completely satisfactory answer. Treatment of composition and concurrency, the inflexibility of the temporal quantifiers, the ad hoc nature of the formal semantics were among the issues that made it difficult to get a handle on the complexity of actual operating system design. Much of the methods used here and in [14] were present in a 1991 paper [13], which included Moore machines, automata products for composition and concurrency and even a modal primitive recursive specification language. But that work was in the tradition of computational formal logic and this is not. When computer scientists started to try to develop mathematical methods of specifying programs, they often adopted methods from formal logic and other axiomatic methods perhaps because of the role of formal languages in programming languages and pattern matching and because of the influence of metamathematics on computability studies. Whatever the merits of that approach, discarding the apparatus of formal logic has allowed solution of a number of problems that blocked progress in the earlier method and greatly simplified specifications.
The quote from Von Neumann that starts this paper is from a, not surprisingly, enormously perceptive paper that among other things previews algorithmic complexity theory. The full paragraph in which that quote is found has an interesting critique which I have tried to learn from as I returned to this work after a long absence.
@@ Appendix A Appendix: Sequence utilities
Appendix B Appendix: Sequential functions
This is a short summary of results from [14]. The standard presentation of a Moore machine is $(A,X,S,s_{0},\delta,\lambda)$ where $s_{0}\in S$ is the start state, $\delta:S\times A\to S$ and $\lambda:S\to X$. Primitive recursive sequential functions are reformulations to make the state set implicit and the recursion explicit.
$D(\epsilon)=c,\ D(w\cdot a)=g(D(w),a)\text{ and }M(w)=L(D(w))$
Such a map can be transformed into a conventional (possibly infinite state) Moore machine by letting the state set be the set $\{D(w):w\in A^{*}\}$.
A sequential function is any map $f:A^{*}\to X$ where $A$ is a set of events, $A^{*}$ are the finite sequences over $A$ and $X$ is the set of outputs. Sequential functions define possibly infinite state Moore machines via the usual congruence on sequences.
The primitive recursive sequence functions are a proper subset with some useful properties including $f(w)$ must be defined if $f(w\cdot a)$ is defined. The finite state sequential functions are a proper subset of the primitive recursive sequential functions that correspond to finite state Moore machines. Specifications of real devices and programs tend to use primitive recursive and finite state primitive recursive sequential functions without any special attention because properties of real systems are constructive and real systems are finite state machines anyway.
Primitive recursive sequential functions are constructed by a finite number of applications of the following rules.
  1. 1.
    Constant. If $c$ is a constant $f(w)=c$ defines a sequential function.
  2. 2.
    Constant map. If $g$ is a non-state dependent map, $f(w,x)=g(x)$ defines a sequential function.
  3. 3.
    Primitive Recursion on Sequences. If $g$ is a non-state dependent map, $c$ is a constant, $\epsilon$ is the empty sequence
    $f(\epsilon)=c,\ f(w\cdot a)=g(a,f(w)))$
    defines a sequential function. If $g$ has a finite image $f$ is finite state.
  4. 4.
    Simple Composition. If $f_{1}$ is a sequential function, and $h(x)$ is an ordinary map
    $f(w)=h(f_{1}(w))$
    defines a sequential function.
As with the arithmetic primitive recursive functions[9], many additional forms can be defined without changing the class of primitive recursive sequential functions. In particular multiple arguments $f(w,x_{1},\dots x_{n})$ are fine, even with some arguments that are also state dependent $f(w)=(f(v_{1}(w),\dots f_{n}(w)))$. An analog of arithmetic simultaneous recursion allows definition of composite system which correspond to automata products and specify concurrent/parallel systems of machines.
  • General Product
    Given sequential functions $f_{1},\dots f_{n}$, each $f_{i}:A_{i}^{*}\to X_{i}$ and appropriate maps $g_{i}:A\times X\to A_{i}^{*}$ (not sequence dependent) where $X=X_{1}\times\dots X_{n}$,
    $f(w)=(f_{1}(u_{1}(w)),\dots f_{n}(u_{n}(w)))$
    $u_{i}(\epsilon)=\epsilon,\text{ and }u_{i}(w\cdot a)=u_{i}(w)\mathop{\mathtt{concat}}g(a,i,v(w))$
As shown in [14]: If each of the $f_{i}$ is primitive recursive, then $f$ is primitive recursive. If each $f_{i}$ is finite state and $A$ is finite then $f$ is finite state.

### v3

@@ 5 Discussion
When Krithivasan Ramamritham gave me temporal logic papers to read[14] at the start of my graduate studies in the 1980s, I had just finished working on a commercial distributed operating system[2] and was frustrated by how often developers found design errors only during test after writing enormous quantities of code. Temporal logic seemed like a solution, but I was not unique in concluding is was not a satisfactory answer after attempting to apply it to real systems. Lamport and Abadi argued (in 1994) that ”It has long been known that the methods originally developed for shared memory multiprocessing apply equally well to distributed systems” (p. 1543) but it’s not clear that they do apply well at all. Much of the complexity of the Abadi and Lamport paper comes from the assumption that there is a “fair scheduler” somewhere deciding which nondeterministic action to activate – for example, to advance the time variable.
This view is reflected in the use of infinite sequences of assignment maps (maps assigning values to formal variables ) as formal semantics. Infinite sequences of assignment maps that must be composed by interleaving and where semantics depends a scheduler that exists outside the quantifiers inflexibility of the temporal quantifiers, the ad hoc nature of the formal semantics were among the issues that made it difficult to get a handle
Some temporal logic variants use infinite sequences of assignment maps (maps from propositions or variables to values) as formal semantics. modal and temporal logics can define state machines or use state machines as semantic models. Kripke’s initial modal logic paper[8] makes this clear although he did not use the terminolgy of state machines. Finite state machines, of a kind, can be constructed from logical assertions to serve as models for temporal logic as well[4, 3]:
Temporal logic formulae are interpreted over a given finite state graph, also called a (Kripke) structure, $M$ comprised of a set $S$ of states, a total binary transition relation $R\subseteq S\times S$, and a labelling $L$ of states with atomic facts (propositions) true there. – Edmond Clarke [5]
These are non-deterministic machines of a sort with unlabeled state transitions which may be controlled by logical predicates. States are assignment maps assigning values to formal symbols, and composition is by interleaving and essentially taking the union of the assigment maps assuming the domains are disjoint. The view is still derived from concurrent processes in a timeshared single processor computer system and there is clear model of composition.
Methods used here and in [17] were present in a 1991 paper [16], which included Moore machines, automata products for composition and concurrency and even a modal primitive recursive specification language. But that work was in the tradition of computational formal logic and this is not. When computer scientists started to try to develop mathematical methods of specifying programs, they often adopted methods from formal logic and other axiomatic methods perhaps because of the role of formal languages in programming languages and pattern matching and because of the influence of metamathematics on computability studies. Whatever the merits of that approach, discarding the apparatus of formal logic has allowed solution of a number of problems that blocked progress in the earlier method and greatly simplified specifications.
The quote from Von Neumann that starts this paper is from an enormously perceptive paper (not unusual for the author) that among other things previews algorithmic complexity theory. The full paragraph in which that quote is found has an interesting critique which I have tried to learn from as I returned to this work after a long absence.